# PGS: Pose-Guided Supervision for Mitigating Clothes-Changing in Person Re-Identification

### 31 Dec 2023 v2

@@ 4.3 Ablation Studies
Cross-domain testing: We compared our framework with several methods: SSL [24], ATNet [25], UDAP [26], and the baseline is the SOLIDER [3]. From Table 6, our method outperforms all previous methods in both cross-domain settings, which reflects the robustness and generalization of our model in cross-domain Re-ID. Notably, these results underscore the pivotal role of the pose knowledge in the Re-ID task.
@@ 4.4 Feature Map Visualization
To gain deeper insights into the behavior of our framework, we conducted heatmap visualization, as depicted in Figure 2. This visualization readily highlights that the resulting heatmaps from the three versions (PGS-1, PGS-2, PGS-3) primarily concentrate on the human’s head and various body parts. In simpler terms, the synergy of the human encoder, pose encoder, and the PHP module equips the model with the ability to accurately prioritize the human body, reducing its dependence on clothing variations.
@@ 5 Conclusion
In conclusion, this paper introduces the concept of Pose Guidance Supervision (PGS), which transfers the pose knowledge into the Re-ID model at multiple scales for robust Re-ID in scenarios involving both clothes-consistent and clothes-changing. Our study demonstrates the effectiveness of this approach in enabling the model to learn robust features, as evidenced by its superior performance compared to current state-of-the-art methods. This simple approach is a promising candidate for real-world camera surveillance applications and provides a solid foundation for future studies aimed at building more robust models. We encourage fellow researchers to explore various methods for extracting human representations and biometric traits to elevate the performance of Re-ID models further.

### 01 Jun 2024 v3

@@ 4.3 Ablation Study
Impact of for knowledge transfer: As illustrated in Table 6, the results indicate that the model’s performance first improves and then declines when equals 1.0. This phenomenon occurs because a low limits the model’s ability to adapt to pose knowledge, while setting to 1.0 can make it challenging for the model to optimize both the triplet loss and guide loss simultaneously.
@@ 4.4 Feature Map Visualization
To gain deeper insights into the behavior of our framework, we conducted heatmap visualization, as depicted in Figure 3. This visualization readily highlights that the resulting heatmaps from the three versions (PGDS-1, PGDS-2, PGDS-3) primarily concentrate on the human’s head and various body parts. In simpler terms, the synergy of the human encoder, pose encoder, and the PHP module equips the model with the ability to accurately prioritize the human body, reducing its dependence on clothing variations.
@@ 5 Conclusion
In conclusion, this paper introduces the concept of Pose-Guidance Deep Supervision (PGDS), which transfers the pose knowledge into the main Re-ID model at multiple scales for robust Re-ID in scenarios involving both clothes-consistent and clothes-changing. Our study demonstrates the effectiveness of this approach in enabling the model to learn robust features, as evidenced by its competitive performance compared to current state-of-the-art methods. This simple approach is a promising candidate for real-world camera surveillance applications and provides a solid foundation for future studies to build more robust models. We encourage fellow researchers to explore various methods to improve this framework and address cloth-changing problems through the approach that integrates pose information.
Acknowledgement. Dinh-Hieu Hoang is funded by Vingroup Joint Stock Company and supported by the Domestic Master/ PhD Scholarship Programme of Vingroup Innovation Foundation (VINIF), Vingroup Big Data Institute (VINBIGDATA), code VINIF.2022.ThS.JVN.04. Debesh Jha and Ulas Bagci are supported by the NIH funding: R01-CA246704, R01-CA240639, U01-DK127384-02S1, and U01-CA268808.